# KiDS-Legacy: Redshift distributions and their calibration

### v1

@@
11institutetext: Ruhr University Bochum, Faculty of Physics and Astronomy, Astronomical Institute (AIRUB), German Centre for Cosmological Lensing, 44780 Bochum, Germany 22institutetext: Center for Theoretical Physics, Polish Academy of Sciences, al. Lotników 32/46, 02-668 Warsaw, Poland 33institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, UK. 44institutetext: Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, UK 55institutetext: Department of Physics, University of Oxford, Denys Wilkinson Building, Keble Road, Oxford OX1 3RH, United Kingdom. 66institutetext: Donostia International Physics Center, Manuel Lardizabal Ibilbidea, 4, 20018 Donostia, Gipuzkoa, Spain. 77institutetext: Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 88institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Herschel Building, NE1 7RU, Newcastle-upon-Tyne, UK 99institutetext: Institute for Theoretical Physics, Utrecht University, Princetonplein 5, 3584CC Utrecht, The Netherlands. 1010institutetext: Leiden Observatory, Leiden University, P.O.Box 9513, 2300RA Leiden, The Netherlands $1$institutetext: Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona), Spain 1212institutetext: Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Av. Complutense 40, E-28040 Madrid, Spain 1313institutetext: Institute of Cosmology & Gravitation, Dennis Sciama Building, University of Portsmouth, Portsmouth, PO1 3FX, United Kingdom 1414institutetext: Dipartimento di Fisica e Astronomia ”Augusto Righi” - Alma Mater Studiorum Università di Bologna, via Piero Gobetti 93/2, 40129 Bologna, Italy 1515institutetext: INAF-Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, Via Piero Gobetti 93/3, 40129 Bologna, Italy 1616institutetext: Universität Innsbruck, Institut für Astro- und Teilchenphysik, Technikerstr. 25/8, 6020 Innsbruck, Austria 1717institutetext: The Oskar Klein Centre, Department of Physics, Stockholm University, AlbaNova University Centre, SE-106 91 Stockholm, Sweden 1818institutetext: Imperial Centre for Inference and Cosmology (ICIC), Blackett Laboratory, Imperial College London, Prince Consort Road, London SW7 2AZ, UK 1919institutetext: Zentrum für Astronomie, Universitatät Heidelberg, Philosophenweg 12, D-69120 Heidelberg, Germany; Institute for Theoretical Physics, Philosophenweg 16, D-69120 Heidelberg, Germany 2020institutetext: Istituto Nazionale di Fisica Nucleare (INFN) - Sezione di Bologna, viale Berti Pichat 6/2, I-40127 Bologna, Italy 2121institutetext: INAF - Osservatorio Astronomico di Padova, via dell’Osservatorio 5, 35122 Padova, Italy $2$institutetext: Institute for Particle Physics and Astrophysics, ETH Zürich, Wolfgang-Pauli-Strasse 27, 8093 Zürich, Switzerland 2323institutetext: Institute for Computational Cosmology, Ogden Centre for Fundament Physics - West, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK. 2424institutetext: Centre for Extragalactic Astronomy, Ogden Centre for Fundament Physics - West, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK
@@
Angus H. Wright awright@astro.rub.de11    Hendrik Hildebrandt 11    Jan Luca van den Busch 11    Maciej Bilicki 22    Catherine Heymans 1133    Benjamin Joachimi 44    Constance Mahony 115566    Robert Reischke 1177    Benjamin Stölzner 11    Anna Wittje 11    Marika Asgari 88    Nora Elisa Chisari 991010    Andrej Dvornik 11    Christos Georgiou $1$    Benjamin Giblin 33    Henk Hoekstra 1010    Priyanka Jalan 22    Anjitha John William 22    Shahab Joudaki 12121313    Konrad Kuijken 1010    Giorgio Francesco Lesci 14141515    Shun-Sheng Li 111010    Laila Linke 1616    Arthur Loureiro 17171818    Matteo Maturi 1919    Lauro Moscardini 141415152020    Lucas Porth 77    Mario Radovich 2121    Tilman Tröster $2$    Maximilian von Wietersheim-Kramsta 23232424    Ziang Yan 11    Mijin Yoon 1010    Yun-Hao Zhang 11331010
(Released 12/12/2121)
We present the redshift calibration methodology and bias estimates for the cosmic shear analysis of the the fifth and final data release (DR5) of the Kilo-Degree Survey (KiDS). KiDS-DR5 includes a greatly expanded compilation of calibrating spectra, drawn from $27$ square degrees of dedicated optical and near-IR imaging taken over deep spectroscopic fields. The redshift distribution calibration leverages a range of new methods and updated simulations to produce the most precise $N(z)$ bias estimates used by KiDS to date. Improvements to our colour-based redshift distribution measurement method (SOM) mean that we are able to use many more sources per tomographic bin for our cosmological analyses, and better estimate the representation of our source sample given the available spec-$z$. We validate our colour-based redshift distribution estimates with spectroscopic cross-correlations (CC). We find that improvements to our cross-correlation redshift distribution measurement methods mean that redshift distribution biases estimated between the SOM and CC methods are fully consistent on simulations, and the data calibration is consistent to better than $2\sigma$ in all tomographic bins.
@@ 1 Introduction
Wide-field imaging surveys with large mosaic CCD cameras and broad-band optical and near-infrared (NIR) filters have entered a crucial era where significant fractions of the sky are being surveyed. The current generation called stage-III (Sevilla-Noarbe et al., 2021; Aihara et al., 2022; Wright et al., 2024) covers areas of more than a thousand square degrees and will soon be superseded by stage-IV surveys (Euclid Collaboration: Mellier et al., 2024; Ivezić et al., 2019) covering an order of magnitude larger areas at similar or greater depths. Perhaps the most crucial analysis step for virtually any application of these surveys is to add information about the radial distance of the very large number of objects (typically of order $10^{7}$$10^{9}$) reliably detected in such surveys. In the absence of spectroscopic redshifts for these huge samples of (mostly) galaxies, photometric redshifts (photo-$z$; for a recent review see Newman & Gruen, 2022) based on broad-band multi-colour photometry are used to solve this problem.
@@
The estimation of these broad-band photo-$z$ for faint targets has been surprisingly stable over the past two decades (Hildebrandt et al., 2010). All stage-III surveys base their main scientific analyses still on template-fitting techniques developed more than 20 years ago (e.g. Benítez, 2000). This reflects the maturity of these techniques and their close-to optimal use of information. Until the arrival of large, complete spectroscopic training sets down to the magnitude limits of these wide-field imaging surveys, which would enable highly precise and accurate photometric redshifts estimated via machine-learning techniques, this situation is unlikely to change (Newman et al., 2015).
@@
These photo-$z$ estimates of individual galaxies have well characterised error distributions with typical scatter of a few per cent around the true redshifts and equally a fraction of a few per cent of catastrophic outliers. These numbers have also been essentially unchanged for a long time. The main reason for this perceived stagnation in individual photo-$z$ quality is the fact that this performance is not the limiting factor for the main science driver of such imaging surveys: weak gravitational lensing (WL).
@@
The gravitational lensing effect is integrated along the line-of-sight and – in the case of WL – measured statistically by averaging over shear estimates of very large ensembles of galaxies. As such, a significant improvement in individual galaxy photo-$z$ is not required. Instead, individual galaxy photo-$z$ values are used only to divide the galaxy distribution into relatively broad, so-called tomographic bins (hundreds of Mpc comoving) along the line-of-sight.
@@
It is the ensemble redshift distribution, $N(z)$, that has – rightly – received most attention in WL measurements of the recent past as its accuracy is directly related to the accuracy of the cosmological parameters estimated from WL surveys. The increasing statistical power, hence, comes with a paralleled increase in the required accuracy of these $N(z)$, most importantly expressed by their mean redshifts (Huterer et al., 2006). Higher-order moments of the $N(z)$ are less important for cosmic shear but very relevant for other probes like galaxy clustering McLeod et al. (2017); Reischke (2024). Here, we concentrate on the former and leave the quantification of calibration uncertainties of the width, skewness, etc. of the $N(z)$ to future work.
For the current-generation stage-III surveys, the mean redshifts have to be controlled at the per-cent level (Myles et al., 2021; Rau et al., 2023; Hildebrandt et al., 2021). Any larger bias in the redshifts would lead to a bias in the cosmological conclusions that would rival the statistical uncertainty. Calibration techniques are used to estimate the $N(z)$ and simulations are employed to estimate residual biases, which can be used to re-calibrate the data. The uncertainty in this re-calibration is typically marginalised over in the cosmological inference.
@@
The Kilo-Degree Survey (KiDS Wright et al., 2024) is conducted with OmegaCam mounted at the Cassegrain focus of the ESO VLT Survey Telescope (VST) on Paranal, Chile, and complemented by the VISTA [^1] Visible and Infrared Survey Telescope for Astronomy Kilo Degree Infrared Galaxy Survey (VIKING Edge et al., 2013) observed from a neighbouring mountaintop. Together, these two surveys form a unique 9-band data set covering the near-UV to near-IR with (in terms of depth) well-matched high-resolution images over an area of 1 350similar-toabsent1350$\sim$ 1\,350∼ 1 350 $\text{deg}^{2}$. With this extensive filter coverage, KiDS has the potential of estimating well-controlled photo-$z$ down to its magnitude limit ($r\sim 24$ for a typical WL source at a signal-to-noise ratio of $\sim 10$) and estimating accurate $N(z)$ all the way to $z\la 2$, paving the way for similar multi-camera, optical+NIR efforts with e.g. Euclid.
@@
While individual galaxy photo-$z$ and their quality for the complete KiDS data set are covered in the data release 5 paper (dr5; Wright et al., 2024), here we describe the redshift calibration approach, i.e. the estimation of the $N(z)$, and their characterisation with simulations. This is the final paper in a list of publications that have developed the KiDS redshift calibration strategy (Hildebrandt et al., 2017, 2020, 2021; Wright et al., 2020a; van den Busch et al., 2020, 2022). Similar to previous efforts, we use two complementary techniques to estimate the $N(z)$, one that is colour-based and another one that is position-based. Both of these techniques leverage the power of spectroscopic surveys that overlap with KiDS or the newly compiled KiDZ data set (i.e. the KiDS redshift calibration fields). The kind of spectroscopic surveys used for the two techniques are quite different though, which is highly beneficial for systematic robustness and independence of these methods.
@@
The KiDS data, the KiDZ calibration fields, the calibrating spectroscopic surveys, and the tomographic binning approach are described in Sect. 2. The mock catalogues that mimic these different data sets are introduced in Sect. 3. In Sect. 4, the colour-based calibration technique via a self-organising map (SOM) projection of the 9-dimensional colour space is introduced. This is complemented by a description of the position-based calibration technique, also known as clustering redshifts (or dubbed CC for cross-correlation), in Sect. 5. The performance of these two approaches is evaluated on the simulated mock catalogues and presented in Sect. 6. Results on the KiDS/KiDZ data are shown in Sect. 7, which are further discussed in Sect. 8, before we summarise in Sect. 9.
@@ 2 Data
VST (OmegaCAM) $u$ 3 55035503\,5503 550 $24.26\pm 0.10$ $1.01\pm 0.17$
$g$ 4 77547754\,7754 775 $25.15\pm 0.12$ $0.88\pm 0.15$
$r$ 6 23062306\,2306 230 $25.07\pm 0.14$ $0.70\pm 0.12$
$i_{1}$ 7 63076307\,6307 630 $23.66\pm 0.25$ $0.81\pm 0.18$
$i_{2}$ 7 63076307\,6307 630 $23.73\pm 0.30$ $0.81\pm 0.18$
VISTA (VIRCAM) $Z$ 8 77087708\,7708 770 $23.79\pm 0.20$ $0.90\pm 0.10$
$Y$ 10 2001020010\,20010 200 $23.02\pm 0.19$ $0.86\pm 0.09$
$J$ 12 5201252012\,52012 520 $22.72\pm 0.20$ $0.85\pm 0.07$
$H$ 16 4501645016\,45016 450 $22.27\pm 0.24$ $0.88\pm 0.09$
$K_{\rm s}$ 21 4702147021\,47021 470 $22.02\pm 0.19$ $0.87\pm 0.08$
@@
The fifth data release of KiDS consists of $1347$ $\text{deg}^{2}$ of weak lensing imaging data, and $27$ $\text{deg}^{2}$ of imaging covering deep spectroscopic calibration fields (with $4$ $\text{deg}^{2}$ of overlap). All data are observed with both VST and VISTA, yielding photometry in nine distinct photometric bandpasses (four optical and five near-infrared). Additionally, the entire wide and calibration footprint was observed twice in the $i$-band, yielding two realisations and epochs of the photometry in this band. These realisations are kept separate in our analysis, and are labelled $i_{1}$ and $i_{2}$ for distinction (the impact of the additional $i$-band measurements on our photo-$z$ is shown in W24). This leads to a final data set containing ten photometric bands, which are summarised in Table 1. Sources in these fields are extracted from the VST $r$-band imaging using Source Extractor (Bertin & Arnouts, 1996), within the Astro-WISE analysis environment (Valentijn et al., 2007; Begeman et al., 2013; McFarland et al., 2013), approximately $139$ million unique sources.
@@
All sources in KiDS-dr5 have photometric information measured in all available photometric bands. This photometric information is estimated through a form of matched aperture photometry that ensures consistent flux information is extracted from each source across the ten photometric bands, based on the optical $r$-band, and accounting for variations in the point spread function (PSF) per-band. This forced photometry is performed with the Gaussian aperture and PSF code (GAaP; Kuijken 2008), and details of the implementation of GAaP in the context of KiDS-dr5 can be found in sections 3.6 and 6 of W24.
@@
After measurement of photometric information in all bands, the KiDS-dr5 is masked to include only unique sources that reside in areas of high-quality data in all bands. This masking process is described at length in section 6.4 of W24, and results in 100 744 685100744685100\,744\,685100 744 685 sources drawn from an effective area of $1014.013$ $\text{deg}^{2}$(corresponding to an effective number density of $10.94\leavevmode\nobreak\ {\text{arcmin}^{-2}}$).
@@
The lensing portion of the KiDS-dr5 sample is given the name KiDS-Legacy. As in previous KiDS analyses, the lensing sample contains per-source shape measurements and corresponding shape-measurement confidence weights estimated using the lensfit algorithm (Miller et al., 2007, 2013). These shapes are then calibrated with complex image simulations designed to emulate the properties of the KiDS-Legacy sample as closely as possible. A detailed description of these simulations is given in Li et al. (2022), and they are also summarised here in Sect. 3. The definition of the sample is provided in detail in section 7.2 of W24, and involves a series of cuts in magnitude, colour, neighbour distance on-sky, and shape-measurement quality metrics. Additionally, Wright et al. (submitted) found that masking of areas with higher astrometric noise was required to satisfy their cosmic shear B-mode null tests, leading to an additional masking of the survey footprint. The final KiDS-Legacy lensing sample is defined as the remaining 40 950 6074095060740\,950\,60740 950 607 sources after these selections, drawn from $967.4$ $\text{deg}^{2}$ (corresponding to an effective number density of $8.81\leavevmode\nobreak\ {\text{arcmin}^{-2}}$).

### v2

@@
11institutetext: Ruhr University Bochum, Faculty of Physics and Astronomy, Astronomical Institute (AIRUB), German Centre for Cosmological Lensing, 44780 Bochum, Germany 22institutetext: Center for Theoretical Physics, Polish Academy of Sciences, al. Lotników 32/46, 02-668 Warsaw, Poland 33institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, UK. 44institutetext: Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, UK 55institutetext: Department of Physics, University of Oxford, Denys Wilkinson Building, Keble Road, Oxford OX1 3RH, United Kingdom. 66institutetext: Donostia International Physics Center, Manuel Lardizabal Ibilbidea, 4, 20018 Donostia, Gipuzkoa, Spain. 77institutetext: Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 88institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Herschel Building, NE1 7RU, Newcastle-upon-Tyne, UK 99institutetext: Institute for Theoretical Physics, Utrecht University, Princetonplein 5, 3584CC Utrecht, The Netherlands. 1010institutetext: Leiden Observatory, Leiden University, P.O.Box 9513, 2300RA Leiden, The Netherlands 1111institutetext: Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona), Spain 1212institutetext: Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Av. Complutense 40, E-28040 Madrid, Spain 1313institutetext: Institute of Cosmology & Gravitation, Dennis Sciama Building, University of Portsmouth, Portsmouth, PO1 3FX, United Kingdom 1414institutetext: Dipartimento di Fisica e Astronomia ”Augusto Righi” - Alma Mater Studiorum Università di Bologna, via Piero Gobetti 93/2, 40129 Bologna, Italy 1515institutetext: INAF-Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, Via Piero Gobetti 93/3, 40129 Bologna, Italy 1616institutetext: Universität Innsbruck, Institut für Astro- und Teilchenphysik, Technikerstr. 25/8, 6020 Innsbruck, Austria 1717institutetext: The Oskar Klein Centre, Department of Physics, Stockholm University, AlbaNova University Centre, SE-106 91 Stockholm, Sweden 1818institutetext: Imperial Centre for Inference and Cosmology (ICIC), Blackett Laboratory, Imperial College London, Prince Consort Road, London SW7 2AZ, UK 1919institutetext: Zentrum für Astronomie, Universitatät Heidelberg, Philosophenweg 12, D-69120 Heidelberg, Germany; Institute for Theoretical Physics, Philosophenweg 16, D-69120 Heidelberg, Germany 2020institutetext: Istituto Nazionale di Fisica Nucleare (INFN) - Sezione di Bologna, viale Berti Pichat 6/2, I-40127 Bologna, Italy 2121institutetext: INAF - Osservatorio Astronomico di Padova, via dell’Osservatorio 5, 35122 Padova, Italy 2222institutetext: Institute for Particle Physics and Astrophysics, ETH Zürich, Wolfgang-Pauli-Strasse 27, 8093 Zürich, Switzerland 2323institutetext: Institute for Computational Cosmology, Ogden Centre for Fundament Physics - West, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK. 2424institutetext: Centre for Extragalactic Astronomy, Ogden Centre for Fundament Physics - West, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK
@@
Angus H. Wright awright@astro.rub.de    Hendrik Hildebrandt    Jan Luca van den Busch    Maciej Bilicki    Catherine Heymans    Benjamin Joachimi    Constance Mahony    Robert Reischke    Benjamin Stölzner    Anna Wittje    Marika Asgari    Nora Elisa Chisari    Andrej Dvornik    Christos Georgiou    Benjamin Giblin    Henk Hoekstra    Priyanka Jalan    Anjitha John William    Shahab Joudaki    Konrad Kuijken    Giorgio Francesco Lesci    Shun-Sheng Li    Laila Linke    Arthur Loureiro    Matteo Maturi    Lauro Moscardini    Lucas Porth    Mario Radovich    Tilman Tröster    Maximilian von Wietersheim-Kramsta    Ziang Yan    Mijin Yoon    Yun-Hao Zhang
(Received 31 March 2025 / Accepted 26 August 2025 )
We present the redshift calibration methodology and bias estimates for the cosmic shear analysis of the fifth and final data release (DR5) of the Kilo-Degree Survey (KiDS). KiDS-DR5 includes a greatly expanded compilation of calibrating spectra, drawn from $27$ square degrees of dedicated optical and near-IR imaging taken over deep spectroscopic fields. The redshift distribution calibration leverages a range of new methods and updated simulations to produce the most precise $N(z)$ bias estimates used by KiDS to date. Improvements to our colour-based redshift distribution measurement method using self-organising maps (SOMs) mean that we are able to use many more sources per tomographic bin for our cosmological analyses and better estimate the representation of our source sample given the available spec-$z$. We validated our colour-based redshift distribution estimates with spectroscopic cross-correlations (CCs). We find that improvements to our CC redshift distribution measurement methods mean that redshift distribution biases estimated between the SOM and CC methods are fully consistent on simulations, and the data calibration is consistent to better than $2\sigma$ in all tomographic bins.
@@ 1 Introduction
Wide-field imaging surveys with large mosaic CCD cameras and broadband optical and near-infrared (NIR) filters have entered a crucial era where significant fractions of the sky are currently being surveyed. The current generation called stage-III (Sevilla-Noarbe et al., 2021; Aihara et al., 2022; Wright et al., 2024) covers areas of more than a thousand square degrees and will soon be superseded by stage-IV surveys (Euclid Collaboration et al., 2025; Ivezić et al., 2019) covering an order of magnitude larger areas at similar or greater depths. Perhaps the most crucial analysis step for virtually any application of these surveys is to add information about the radial distance of the very large number of objects (typically of the order of $10^{7}$$10^{9}$) reliably detected in such surveys. In the absence of spectroscopic redshifts for these huge samples of (mostly) galaxies, photometric redshifts (photo-$z$; for a recent review see Newman & Gruen, 2022) based on broadband multi-colour photometry are used to solve this problem.
@@
The estimation of these broadband photo-$z$ for faint targets has been surprisingly stable over the past two decades (Hildebrandt et al., 2010). All stage-III surveys base their main scientific analyses still on template-fitting techniques developed more than 20 years ago (e.g. Benítez, 2000). This reflects the maturity of these techniques and their close-to optimal use of information. Until the arrival of large, complete spectroscopic training sets down to the magnitude limits of these wide-field imaging surveys, which would enable highly precise and accurate photometric redshifts estimated via machine-learning techniques, this situation is unlikely to change (Newman et al., 2015).
@@
These photo-$z$ estimates of individual galaxies have well characterised error distributions with typical scatter of a few per cent around the true redshifts and equally a fraction of a few per cent of catastrophic outliers. These numbers have essentially remained unchanged for a long time. The main reason for this perceived stagnation in individual photo-$z$ quality is the fact that this performance is not the limiting factor for the main science driver of such imaging surveys: weak gravitational lensing (WL).
@@
The gravitational lensing effect is integrated along the line of sight and – in the case of WL – measured statistically by averaging over shear estimates of very large ensembles of galaxies. As such, a significant improvement in individual galaxy photo-$z$ is not required. Instead, individual galaxy photo-$z$ values are used only to divide the galaxy distribution into relatively broad, so-called tomographic bins (hundreds of Mpc comoving) along the line of sight.
@@
It is the ensemble redshift distribution, $N(z)$, that has rightly received most attention in WL measurements taken in the recent past as its accuracy is directly related to that of the cosmological parameters estimated from WL surveys. The increasing statistical power, hence, comes with a paralleled increase in the required accuracy of these $N(z)$, most importantly expressed by their mean redshifts (Huterer et al., 2006). Higher-order moments of the $N(z)$ are less important for cosmic shear but very relevant for other probes like galaxy clustering McLeod et al. (2017); Reischke (2024). Here, we concentrate on the former and leave the quantification of calibration uncertainties on higher-order moments of the $N(z)$, such as width and skewness, to future work. For the current-generation stage-III surveys, the mean redshifts must be controlled at the percent level (Myles et al., 2021; Rau et al., 2023; Hildebrandt et al., 2021). Any larger bias in the redshifts would lead to a bias in the cosmological conclusions that would rival the statistical uncertainty. Calibration techniques are used to estimate the $N(z)$, and simulations are employed to estimate residual biases, which can be used to re-calibrate the data. The uncertainty in this re-calibration is typically marginalised over in the cosmological inference.
@@
The Kilo-Degree Survey (KiDS; Wright et al., 2024) is conducted with OmegaCam mounted at the Cassegrain focus of the European Southern Observatory VLT Survey Telescope (VST) on Paranal, Chile and complemented by the VISTA [^1] Visible and Infrared Survey Telescope for Astronomy Kilo Degree Infrared Galaxy Survey (VIKING; Edge et al., 2013) observed from a neighbouring mountaintop. Together, these two surveys form a unique nine-band dataset covering the near-UV to NIR with (in terms of depth) well-matched high-resolution images over an area of $\sim 1\,350$ $\text{deg}^{2}$. With this extensive filter coverage, KiDS has the potential of estimating well-controlled photo-$z$ down to its magnitude limit ($r\sim 24$ for a typical WL source at a S/N of $\sim 10$) and estimating accurate $N(z)$ all the way to $z\la 2$, paving the way for similar multi-camera, optical+NIR efforts with, for example, Euclid.
@@
While individual galaxy photo-$z$ and their quality for the complete KiDS dataset are covered in the Data Release 5 (DR5) paper (Wright et al., 2024), here we describe the redshift calibration approach; that is, the estimation of the $N(z)$, and their characterisation with simulations. This is the final paper in a list of publications that have developed the KiDS redshift calibration strategy (Hildebrandt et al., 2017, 2020, 2021; Wright et al., 2020a; van den Busch et al., 2020, 2022). Similar to previous efforts, we used two complementary techniques to estimate the $N(z)$: one that is colour-based and another that is position-based. Both of these techniques leverage the power of spectroscopic surveys that overlap with KiDS or the newly compiled KiDZ dataset (i.e. the KiDS redshift calibration fields). The kinds of spectroscopic surveys used for the two techniques are quite different, though, which is highly beneficial for systematic robustness and independence of these methods.
@@
The KiDS data, the KiDZ calibration fields, the calibrating spectroscopic surveys, and the tomographic binning approach are described in Sect. 2. The mock catalogues that mimic these different datasets are introduced in Sect. 3. In Sect. 4, the colour-based calibration technique via a self-organising map (SOM) projection of the 9D colour space is introduced. This is complemented by a description of the position-based calibration technique, also known as clustering redshifts (or dubbed CC for cross-correlation), in Sect. 5. The performance of these two approaches was evaluated on the simulated mock catalogues and is presented in Sect. 6. Results on the KiDS and KiDZ data are shown in Sect. 7, which are further discussed in Sect. 8 before we summarise in Sect. 9.
@@ 2 Data
VST (OmegaCAM) $u$ 3 5503\,550 $24.26\pm 0.10$ $1.01\pm 0.17$
$g$ 4 7754\,775 $25.15\pm 0.12$ $0.88\pm 0.15$
$r$ 6 2306\,230 $25.07\pm 0.14$ $0.70\pm 0.12$
$i_{1}$ 7 6307\,630 $23.66\pm 0.25$ $0.81\pm 0.18$
$i_{2}$ 7 6307\,630 $23.73\pm 0.30$ $0.81\pm 0.18$
VISTA (VIRCAM) $Z$ 8 7708\,770 $23.79\pm 0.20$ $0.90\pm 0.10$
$Y$ 10 20010\,200 $23.02\pm 0.19$ $0.86\pm 0.09$
$J$ 12 52012\,520 $22.72\pm 0.20$ $0.85\pm 0.07$
$H$ 16 45016\,450 $22.27\pm 0.24$ $0.88\pm 0.09$
$K_{\rm s}$ 21 47021\,470 $22.02\pm 0.19$ $0.87\pm 0.08$
@@
The fifth data release of KiDS consists of $1347$ $\text{deg}^{2}$ of weak lensing imaging data, and $27$ $\text{deg}^{2}$ of imaging covering deep spectroscopic calibration fields (with 44 $\text{deg}^{2}$ of overlap). All data were observed with both VST and VISTA, yielding photometry in nine distinct photometric bandpasses (four optical and five NIR). Additionally, the entire wide and calibration footprint was observed twice in the $i$ band, yielding two realisations and epochs of the photometry in this band. These realisations are kept separate in our analysis and are labelled $i_{1}$ and $i_{2}$ for distinction (the impact of the additional $i$-band measurements on our photo-$z$ is shown in W24). This leads to a final dataset containing ten photometric bands, which are summarised in Table 1. Sources in these fields were extracted from the VST $r$-band imaging using Source Extractor (Bertin & Arnouts, 1996), within the Astro-WISE analysis environment (Valentijn et al., 2007; Begeman et al., 2013; McFarland et al., 2013), yeilding approximately $139$ million unique sources.
@@
All sources in KiDS-dr5 have photometric information measured in all available photometric bands. This photometric information was estimated through a form of matched aperture photometry that ensures consistent flux information is extracted from each source across the ten photometric bands, based on the optical $r$-band, and accounting for variations in the point spread function (PSF) per band. This forced photometry was performed with the Gaussian aperture and PSF code (GAaP; Kuijken 2008), and details of the implementation of GAaP in the context of KiDS-dr5 can be found in Sections 3.6 and 6 of W24.
@@
After measurement of photometric information in all bands, the KiDS-dr5 sample was masked to include only unique sources that reside in areas of high-quality data in all bands. This masking process is described at length in Section 6.4 of W24 and results in 100 744 685100\,744\,685 sources drawn from an effective area of $1014.013$ $\text{deg}^{2}$(corresponding to an effective number density of $10.94\penalty 10000\ {\text{arcmin}^{-2}}$).
@@
The lensing portion of the KiDS-dr5 sample was given the name KiDS-Legacy. As in previous KiDS analyses, the lensing sample contains per-source shape measurements and corresponding shape-measurement confidence weights estimated using the lensfit algorithm (Miller et al., 2007, 2013). These shapes were then calibrated with complex image simulations designed to emulate the properties of the KiDS-Legacy sample as closely as possible. A detailed description of these simulations is given in Li et al. (2023), and they are also summarised here in Sect. 3. The definition of the sample is provided in detail in Section 7.2 of W24, and involves a series of cuts in magnitude, colour, neighbour distance on-sky, and shape-measurement quality metrics. Additionally, Wright et al. (2025) found that masking of areas with higher astrometric noise was required to satisfy their cosmic shear B-mode null tests, leading to an additional masking of the survey footprint. The final KiDS-Legacy lensing sample is defined as the remaining 40 950 60740\,950\,607 sources after these selections, drawn from $967.4$ $\text{deg}^{2}$ (corresponding to an effective number density of $8.81\penalty 10000\ {\text{arcmin}^{-2}}$).